UT 12 Sep 2021 - Restart Night #9




Night Staff:

OSA: D.G. Huerta

ISp: Mtn Staff

ISA/Observers: J.Power

Technical Wizard: J. Hill (Tucson)

Mtn Manager: M. Wagner

Sunset: 18:35

Twlight Ends: 19:54

Twilght Starts: 04:38

Sunrise: 05:57

Starting Config: LUCI-Bino (LUCI-1 authorized)

Night Log:

  • Before Opening
    • reset MODS2 slitmasks - "slitmask reset" in MODS-2 GUI
    • Put in tipt tilt and IE CA for field aberration SX - pending email from J.Hill


Proposed Plan of Action:

  • LUCI:
    • DONE LUCI-1 Dual BLR target - twilight until 30deg limit (Rothberg)
      • Take Telluric first as soon as AGW-1 can wave-front sense
  • MODS - if scripts are prepared in time - switch to MODS for these targets
    • MODS (Wagner)
      • DONE 18:57:30 +16:53:39 - 30deg limit at 00:15hr
      • DONE 17:50:13 -06:42:28 - 30deg limit at 22hr
  • If Transforms ready proceed to appropriate instrument to collect Transform Confirmatory data - LUCI first then MODS (or whichever is done) - highest priority
  • LUCI:
    • DONE Confirmatory Field Aberration take as LFBG - 1-1.5hr
      • RFBG - ok as-is - take data anyways
        • Fields:
          • StoneP1 22:37:45.0 -01:13:40.45
          • BS9188 23:30:33.4 +38:18:57.0
          • BS9107 03:32:03.0 +37:20:40.0
    • LUCI-1 NS Longslit
      • Telluric ahead of time
      • Conditions - Any
      • Comet 67 CG - scripts based on UT times - earlies is 01hr local time
        • Enter '90000693' in the NSIGUI (this is the JPL ID Horizon and NSIGUI will work with correctly)
        • IF NSIGUI issues persist, there is a 90000693 txt file in /lbt/ephemeris on ROBS for every UT data through end of restart that can be used
        • dither along the slit +/-45" zJHKspec G200
    • AGW-1/2 Transform Confirmatory Data - 1.5hr - Highest Priority
      • Fields:
      • StoneO 21 37 31 +00 31 55 Stone_O_m12.list
      • StoneP 22 35 49 +01 31 33 Stone_P1.list
      • StoneA 00 54 02 -00 32 34 Stone_A3.list
    • Photometric Standards (clear any seeing)
  • Configure to MODS-Bino
    • Confirmatory AGW Transform - 1.5hr - Highest Priority
    • Spectrophotometric stars (if clear any seeing)
    • Standard Stars (if clear any seeing)
    • Pointing Model (poor weather and poor seeing)
  • Mixed Mode Test - if time
    • LBC-Red & MODS-1
  • LBC:
    • Photometric Standards (clear any seeing)
    • Spectrophotometric stars (clear any seeing)
    • Pointing Model LBC (poor seeing, clouds)
      • Turn on PCS logging
      • Turn on TMS passive mode
      • Need to collimate at each star - or - collimate every few stars
      • Chip 2 only exposure
      • File Location:
        • /home/lbto/Calib_OBs/PtStars_10m_OBs/VV

        • PtStarrs: /home/lbto/POINTING/PtStars_10m/PtStars_10m.dat
    • If conditions too poor for other stuff - Pointing Model in Active mode too

  • Poor Weather (if pointing models are exhausted)
    • LUCI-1 local mergers
    • Service_Observing/Restart_2021/LUCI/PoorWeather/COBandhead_Mergers_V17.xml - OT file
    • finder charts in sub-directory (but they are so large you can't miss 'em)


Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS
    • MODS-1 - issues with M1RC headers & filenumbers short 1 digit
      • AGW-5 - available
    • MODS-2 - available
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • AGW-2 available
  • PEPSI - available
    • AGW-7/8 available


Closed Dome Activities

  • MODS-1 Full set of Cals
    • Biases
      • 1x1
      • 1x2
      • 2x2
    • Grating
      • Dual
      • Red
      • Blue
    • Imaging
      • Dual
      • Red
      • Blue
    • Prism
      • Dual
      • Red
      • Blue
  • MODS-2 Cals
    • Bias
      • 1x2
      • 2x2
    • Prism
      • Dual
      • Red
      • Blue
  • LUCI-1 Cals
    • Imaging N375 Flats
      • Narrow, medium-band filters
      • Broad zJHKKs and HeI completed
    • N1.8 G210 CWL=2.3 K-band 0.5" cals
    • N1.8 G200 CWL=1.93 HKspec 0.5" cals
  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout (alternate to on-sky method)
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)

Night Summary

For the night of: Sat, 2021 Sep 11 ---> Sun, 2021 Sep 12
Local midnight = 2021 Sep 12,  7 hr UT, or JD 2459469.792
Local Mean Sidereal Time at midnight =  23 06 29.2

Sunset ( 1000 m horizon):   18 35 MST; Sunrise:   5 57 MST
Evening twilight:  19 54 MST;  LMST at evening twilight:  19 00
Morning twilight:   4 38 MST;  LMST at morning twilight:   3 45
12-degr twilight: 19 24 MST -->  5 08 MST; night center:   0 16 MST

Moonset :  21 50 MST   
Moon at civil midnight: illuminated fraction 0.318
1.6 days before first quarter, RA and dec:  15 44 34, -19 51.9

The sun is down for 11.4 hr;  8.7 hr from eve->morn 18 deg twilight.
 6.8 dark hours after moonset and before beginning of twilight.

MODS and LUCI have been brought out and checked out.   John put in tipt tilt and IE CA for field aberration SX

Investigated the slitmask faults, moving the mask from one end to the oteher through the cabinet and executing a slitmask reset. Continued to see the faults after the slitmask reset. We need to have some boots on teh ground to listen and watch as masks are moved.
[2021-09-12T00:00:24.076] MC2>M2.IE SLITMASK 24
[2021-09-12T00:00:25.332] M2.IE>MC2 STATUS: Retracting 2
[2021-09-12T00:00:46.477] M2.IE>MC2 ERROR: SLITMASK SLITMASK=2 Move Fault, position at end of move 2 but requested position 24
[2021-09-12T00:01:23.873] MC2>M2.IE slitmask 24
[2021-09-12T00:01:33.937] M2.IE>MC2 DONE: SLITMASK SLITMASK=24 MASKPOS=IN MASKNAME='ID531713'
[2021-09-12T00:02:18.072] MC2>M2.IE SLITMASK 2
[2021-09-12T00:02:19.348] M2.IE>MC2 STATUS: Retracting 24
[2021-09-12T00:02:40.845] M2.IE>MC2 ERROR: SLITMASK SLITMASK=1 Move Fault, position at end of move 1 but requested position 2
[2021-09-12T00:04:00.167] MC2>M2.IE slitmask 2
[2021-09-12T00:04:10.203] M2.IE>MC2 DONE: SLITMASK SLITMASK=2 MASKPOS=IN MASKNAME='SieveMask'
[2021-09-12T00:04:28.171] MC2>M2.IE SLITMASK 3
[2021-09-12T00:04:29.453] M2.IE>MC2 STATUS: Retracting 2
[2021-09-12T00:04:41.077] M2.IE>MC2 ERROR: SLITMASK SLITMASK=2 Move Fault, position at end of move 2 but requested position 3

Some cloud overhead at sunset. David has put the 26 degree elevation limit in place so that we can follow the first target into the floor.

Dual BLR

2:02 Preset to HIP71172, telluric for Dual BLR Science. Still too bright. And we now have thicker clouds at the pointing.

2:15 Remade with 11th mag guide star. Sending preset. Much better. Offset X-shift 10.0495, Y-Shift -0.8034

2:23 Starting telluric science. Seeing 1.1" on the dimm. luci1.20210912.0016-0017

2:27 Slewing to primary science for Dual BLR. Still some clouds overhead. 0.8-1.0" on the guider. X-Shift -0.4510, Y-shift -0.0649, Angle 0.6048. Unable to see secondary source. Heavy cloud or is it a bad alignment. The secondary source is the top priority so redoing alignment.

2:57 Verified alignment and starting science. We have a bit of a break in the cloud to the west now. Seeing 0.9" on DIMM and and on teh guiders. Elevation 36. 36 mnutes until this object hits elevation limits. luci1.20210912.0023, 26-44.

3:59 Target about to hit 26 degree elevation limit. Stopping and Reconfiguring to binocular MODS

RS Oph

4:16 David is executing a pointing check near RS Oph

4:30 Slewing to RS Oph. PA set to 45. The acquistioing images on the guiders look very different between teh 2 sides so I think we picked up the wrong guide star on one side. We will see.

4:35 Right side is about 70" off of desired pointing. David is doing a pointing correction and check.

4:43 Resending. PA 46. Pointing looks more consistent with what we would expect. So correct guide star this time. But when ran script something happened and claims to be in stow in header for all images. Issue with abort previously executed?

MODS1R> mods1r.20210912.0005.fits OBJECT RS Oph Acquisition r_sdss LS5x60x0.3 STOW * mask in place, so why stow?
MODS2R> mods2r.20210912.0005.fits OBJECT RS Oph Acquisition r_sdss LS5x60x0.3 STOW * mask in place? Mask not in place..
MODS1R> mods1r.20210912.0006.fits OBJECT RS Oph Acquisition r_sdss LS5x60x0.3 STOW
MODS2R> mods2r.20210912.0006.fits OBJECT RS Oph Acquisition r_sdss LS5x60x0.3 STOW

4:49 Rerunning acqBinoMODS. PA 46. We have the 26 degree elevation limit so can follow this for a bit longer.

4:48 Science. Not saturated in brightest emission lines so will keep exposure time. Partly cloudy. 1.05" on the guiders.
mods1r.20210912.0011-15
mods2r.20210912.0010-14
mods[1/2]b.20210912.0003-7

v1674

5:14 Slewing to v1674 with PA 58. Partly cloudy, although better than earlier. Seeing 0.6-0.8" on the guiders. The finder chart was not very good. No indication of what was north, or scale. The target was pretty obvious, and verified with Olga's modsView image.

5:25 starting science on v1674. Thin cirrus overhead. Seeing 0.8-1.0"

mods1r.20210912.0019-21
mods2r.20210912.0018-20
mods[1/2]b.20210912.0008-10

Feige110

6:04 Slewing to Feige 110. Light cirrus, thinning overhead. Seeing 1."

6:11 Starting science on Feige 110.

mods1r.20210912.0024-26
mods2r.20210912.0023-25
mods[1/2]b.20210912.0011-13

6:23 Reconfiguring to Binocular LUCI for the field aberration confirmation data collection. Closing the MODS hatch and leaving in a configuration for cals in the morning.

Field Aberration Confirmation Data

Field selected BS9188 23:30:33.4 +38:18:57.0

6:38 Doing a pointing and collimation check near target. Conditions mostly clear. Seeing 0.75" on the guider. DX stuggled to collimate.

7:00 Starting field_collect on BS9188. Seeing 0.6-0.6". Mostly clear skies.

7:20 DX is stuggling to collimate and maintain collimation. Seeing 0.6"

7:38. Seeing deteriorating, highly variable. 0.7-1.1". Making it difficult to collimate.

7:47 SX is struggling to collimate

8:05 DX has completed the data collection. SX has one more star in the list. SX has struggled more with collimation.

8:08 Seeing has dropped back down to 0.5-0.6"

Field Aberration Results

Data lives in /home/lbto/data/20210912/sx_field/ and dx_field/

Summary by JMH: Seeing was variable, so the data is useful but not fantastic. The new Tilt value is solidly confirmed. The new Tip value seemed to be overshot, so I have removed it. The net M2 corrections for this week are thus: Tip=0 Tilt=-25 arcsec.

X=+1.277 Y=-0.0 Z=0 RX=0.0 RY=+245.5

These values have been moved out of "Global Offset" and into "Offset" to keep the "Global Offset" buffer clear for short-term changes.

David: The pointing for all the Gregorian Focal Stations will be adjusted by deltaIE = +15 deltaCA = 0 compared to last night = 20210912 UT. Or, the pointing for all the Gregorian Focal Stations will be adjusted by deltaIE = 0 deltaCA = +25 compared to 2 nights ago (before 20210912 UT).

All SX Gregorian Focal Stations have been updated in the PSF config files SXSM*Collimation.dat - including those for ARGOS.

SX Results
 => Field Data Results
      Motion of the secondary needed to remove bi-nodal astigmatism
  
                  BS9188_field
              From residual Focus Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -14.6", Delta IE =  14.6"
         Tilt =  29.1", Delta CA = -29.1"
  
  
 => Field Data Results
             Motion of the secondary needed to remove bi-nodal astigmatism
  
                  BS9188_field
              From Total Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   0.3", Delta IE =  -0.3"
         Tilt =   0.9", Delta CA =  -0.9"
  
  
 => Field Data Results
             Measured bi-nodal astigmatism
  
                  BS9188_field
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -10.2", Delta IE =  10.2"
         Tilt =   7.1", Delta CA =  -7.1"
  
  
                  BS9188_field
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -18.2", Delta IE =  18.2"
         Tilt =   5.4", Delta CA =  -5.4"
  
  
                  BS9188_field
              From both Astigmatisms Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -14.6", Delta IE =  14.6"
         Tilt =  -1.8", Delta CA =   1.8"

DX results
 => Field Data Results
      Motion of the secondary needed to remove bi-nodal astigmatism
  
                  BS9188_field
              From residual Focus Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -11.4", Delta IE =  11.4"
         Tilt =  -5.7", Delta CA =   5.7"
  
  
 => Field Data Results
             Motion of the secondary needed to remove bi-nodal astigmatism
  
                  BS9188_field
              From Total Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   3.9", Delta IE =  -3.9"
         Tilt =   4.3", Delta CA =  -4.3"
  
  
 => Field Data Results
             Measured bi-nodal astigmatism
  
                  BS9188_field
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  -7.9", Delta IE =   7.9"
         Tilt = -26.4", Delta CA =  26.4"
  
  
                  BS9188_field
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -23.0", Delta IE =  23.0"
         Tilt =  -2.1", Delta CA =   2.1"
  
  
                  BS9188_field
              From both Astigmatisms Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -11.0", Delta IE =  11.0"
         Tilt = -23.3", Delta CA =  23.3"

Longslit Comet 67 CG

8:13 David is reauthorizing as Monocular LUCI1, setting NSI GUI for 90000693.

8:22 The scripts appeared to fail, witht he telescope component showing red,. But while investigating the preset just started to go with no interference from me.

8:25 No guide star found. David is executing a pointing check. He noted some oddness in the IIF log so is reauthorizing.

8:35 StartingLUCI longslit acquisition. There is a problem with the ephemrides being genrated with teh NSIGUI

Those generated today are different from those generated last week.

From today: 2021-Sep-12 08:30:00.000 2459469.854166667 03 42 53.752149 +16 22 50.24397 131.7730 43.05129 12.061 n.a. 99.4193 44.8237
From last week: 2021-Sep-12 08:30:00.000 2459469.854166667 03 42 34.193261 +16 20 55.75006 131.6353 43.07107 11.761 14.809 99.5133 44.8755

jpl is down. Using the rates generated today the target is not in the field. Using last weeks rates we were able to identify the target and the RA rates appear different.

NS Uranus

9:19 Sending track preset to Uranus to see if the ephemeris issue is limited to the comet. JPL web query ok for Uranus but not comet. Uranus lands nice and centered.

9:29 Executing a longslit test on Uranus

9:38 Science started.

10:16 Telluric for Uranus. Offsets X: -10.2474, Y -0.4971

Mixed Mode Test

10:27 Slewing to zenith for a reconfig to MODS1/LBCR

monocular LBCR for collimation with LBCR. ran ptautoadjust. TMS is passively on on LBCR. Edited scripts for Q0636+860 for LBCR to contain the <syncoffset> command.

Missed the --bino when sending the MODS acquisition the first time which caused the preset to fail. Went step by step through and caught my mistake.

11:43 running mods acq and LBCR science


MODS offsets -1.002,10.989

** ERROR: SLITMASK Unknown position ID 'ID521532', must be one of [DarkMask,SieveMask,Imaging,LS5x60x0.3,LS5x60x0.6,LS5x60x0.8,LS5x60x1.0,LS5x60x1.2,LS60x5,PinholeSlit,PointingMask,LS5x60x2.4,Empty,Empty,Empty,Empty,Empty,Empty,Empty,ID571722,ID535231,ID532732,ID531956,ID531713]
Science script has a mask in it. Regenerated with 0.8" which was aligned.

Seeing 1.3", light cirrus

mods1r.20210912.0030-31

12:11 Aborting scripts, getting bright. Closing up. Andrew has turned off lasers

12:28. David is reconfiguring LBC's out of the beam.

Calibrations

MODS 8K biases, dual grating arcs, 0.8" dual grating flats, 1.0" dual grating flats, 0.3" dual grating flats, dual slitless flats
MODS red grating arcs, mods red grating 0.8" flats, red only slitless

LUCI darks, Luci1 0.5" G200 HK arcs & flats, LUCI1 0.5" G200 zJ arcs and flats, LUCI1 0.5" G210 K arcs and flats, LUCI1 1.0" G200 HK arcs and flats
LBC 25 Biases

-- BarryRothberg - 11 Sep 2021
Topic revision: r7 - 12 Sep 2021, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback